OpenStudy (anonymous):

algebraically determine whethere the fuction f(x)=2X times the square root 3+x^4 is even, odd, or neither.

OpenStudy (ybarrap):

f(x) = 2x(3 + x^4) = 6x + 2x^5 f(-x) = -6x - 2x^5 = -(6x + 2x^5) = -2x(3 + x^4) since f(-x) = -f(x), f(x) is odd

OpenStudy (anonymous):

\[f(x)= 2x \sqrt{3+x^4}\]

OpenStudy (anonymous):

f(x) = 2x * sqrt(3+x^4 ) ===> f(-x)= -2x*sqrt(3+x^4 ) ==> -f(x) = -2x *sqrt(3+x^4 ) ==> f(-x)=-f(x) is odd

OpenStudy (anonymous):

can u write that out im confused

OpenStudy (anonymous):

thanks